Following is a list of domestic output and national income figures for a given year. All figures are in billions. Calculate GDP by both the expenditure and income methods. The answers derived by each approach should be the same.


Personal consumption expenditures $120
Capital consumption allowances (depreciation)20
Interest and miscellaneous investment income 10
Net income of farms and unincorporated business 17
Net exports+13
Profits of corporation and government enterprises before taxes 42
Wages, salaries, and supplementary labour income 113
Indirect business taxes (less subsidies)11
Government current purchases of goods and services 40
Net investment (net capital formation) 30
Taxes less subsidies on factors of production 10
1
Expert's answer
2021-09-13T11:21:51-0400

According to the given information, GDP using expenditure method would be:

=Personal consumption expenditures+Net exports+Government current

 purchases of goods and services +Net investment (net capital formation)

"=120+13+40+30\\

=203"


similarly, GDP using income method would be:

=Capital consumption allowances (depreciation)+Interest and miscellaneous investment income+Net income of farms and unincorporated business+

     Profits of corporation and government enterprises before taxes+Wages,salaries, and supplementary labour income+Indirect business taxes

 (less subsidies)−Taxes less subsidies on factors of production 

"=20+10+17+42+113+11−10\\

=203"